OpenStudy (anonymous):

More Trigonometric Equations! :) .. :'( Solve for (0,2pi): sin2xsinx+cos2x+cosx=1 if you could explain the steps too that would be AMAZING! :)

OpenStudy (anonymous):

please.. im struggling i wasn't in class for this lesson..

OpenStudy (anonymous):

Do your know the double angle formula

OpenStudy (anonymous):

yea i know all the identities

OpenStudy (anonymous):

why dont u substitute for sin 2x and cos 2x

OpenStudy (anonymous):

i know you chance the sin2x and cos2x with 2sinxcosx and cos^2x-sin^2x but what do you do then?

OpenStudy (anonymous):

sin2xsinx+cos2x+cos=1 2sinxcoxsinx + 1-2sin^2 x+cos x =1 ?

OpenStudy (anonymous):

uhhh wait up ..one second lemme do that.

zepdrix (zepdrix):

Oh boy this one is a doozy :u

OpenStudy (anonymous):

tell me about it :/

OpenStudy (anonymous):

wait how did you get that @thushananth01 ?

zepdrix (zepdrix):

He went offline :c sec I can help.

OpenStudy (anonymous):

okay thank you! :)

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{\sin2x} \sin x + \color{orangered}{\cos2x} + \cos x =1\]So we need to apply our Double Angle Identities to these two colored parts, yes?

OpenStudy (anonymous):

yea 2sinxcosx and cos^2 x-sin^2 x

zepdrix (zepdrix):

Ok good. Do you remember the other forms for the Cosine Double Angle? Because we want to use this one:\[\Large\rm \color{royalblue}{\sin x \cos x} \sin x + \color{orangered}{1-2\cos^2x} + \cos x =1\]

zepdrix (zepdrix):

(In orange)

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

i wrote the question wrong there is not addition sign between cos2x and cosx

OpenStudy (anonymous):

im sorry! :/

zepdrix (zepdrix):

Multiply?

OpenStudy (anonymous):

yeap

zepdrix (zepdrix):

Mmm ok sec, let's start over. \[\Large\rm \color{royalblue}{\sin2x} \sin x + \color{orangered}{\cos2x}\cos x =1\]

zepdrix (zepdrix):

I posted my Sine Double Angle wrong >.< my bad.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

its okay! i saw lol :)

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{2\sin x \cos x} \sin x + \color{orangered}{(\sin^2x-\cos^2x)}\cos x =1\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

do i cancel out the sinx with the 2sinx?

zepdrix (zepdrix):

No, you multiply them together, giving you a sin^2x, yes?\[\Large\rm 2\sin^2 x \cos x + \color{orangered}{(\sin^2x-\cos^2x)}\cos x =1\]

OpenStudy (anonymous):

yeap that looks right

zepdrix (zepdrix):

Umm this is a little tricky. Let's seeeeee.... I guess we'll write everything in terms of cosine. So rewrite our sin^2x as (1-cos^2x) in both places.

zepdrix (zepdrix):

\[\Large\rm 2(1-\cos^2x) \cos x + \color{orangered}{(1-\cos^2x-\cos^2x)}\cos x =1\]

zepdrix (zepdrix):

I know I know I'm stealing all the fun >.< lol sorry I like my equation tool hehe

OpenStudy (anonymous):

lol its all good :)

zepdrix (zepdrix):

So then we combine like terms, expand everything out, then combine like terms again -_-

zepdrix (zepdrix):

\[\Large\rm 2(1-\cos^2x) \cos x + \color{orangered}{(1-2\cos^2x)}\cos x =1\]\[\Large\rm 2\cos x-2\cos^3x+\cos x-2\cos^3x =1\]\[\Large\rm 2\cos x-2\cos^3x+\cos x-2\cos^3x -1=0\]Oh boy this isn't working out nicely.. hmmm thinking +_+

OpenStudy (anonymous):

jeez i didnt think it would take this long for this problem. maybe is was ment to be written in terms of sine?

zepdrix (zepdrix):

Probably not, since there is a cosx multiplying each term. There wouldn't be a nice way to deal with those.

OpenStudy (anonymous):

oh do you think you can get the answer? :o

zepdrix (zepdrix):

Yah I'll throw it into Wolfram, that should help me see what's going on here >.<

OpenStudy (anonymous):

okay :) cool

zepdrix (zepdrix):

Ohhh I'm so dumb. Yah I forgot to "rethink the problem" when we made the adjustment. This is just simply the Sine Additive Identity.\[\Large\rm \sin(a+b)=\sin a \cos b + \sin b \cos a\]That look familiar?

OpenStudy (anonymous):

ohhhhhh snap ! i feel even stupider lol

zepdrix (zepdrix):

Do you see it popping out at you? :O

OpenStudy (anonymous):

im going to erase my work and start over :) can i tell you if i need help? :/

zepdrix (zepdrix):

Sure :U I'm hittin the sack soon, but I'll be on for another 20 prolly.

OpenStudy (anonymous):

okay ill try to work fast

zepdrix (zepdrix):

Woops woops woops. We have the form sin2xsinx + cos2xcosx I should have posted the Cosine Additive Identity, not the Sine. \[\Large\rm \cos(a-b)=\cos a \cos b+\sin a \sin b\]

OpenStudy (anonymous):

are you sure its not the cosine sum identity?

zepdrix (zepdrix):

>.<

OpenStudy (anonymous):

ohhh okay lolol

zepdrix (zepdrix):

And it looks like it's the difference, not the addition one, right? Did I get that correct? Yah I think my sign is ok.

OpenStudy (anonymous):

yea your right

OpenStudy (anonymous):

hey im stuck >.< @zepdrix

OpenStudy (anonymous):

i dont see it poping out .. -_-

zepdrix (zepdrix):

So our formula is: \[\Large\rm \cos\left(\color{royalblue}{a}-\color{orangered}{b}\right)=\cos\color{royalblue}{a}\cos\color{orangered}{b}+\sin\color{royalblue}{a}\sin\color{orangered}{b}\]And our problem is:\[\Large\rm \cos\color{royalblue}{2x}\cos\color{orangered}{x}+\sin\color{royalblue}{2x}\sin\color{orangered}{x}=1\]yes? Is it popping yet? :U

OpenStudy (anonymous):

isnt x just equal 1?

zepdrix (zepdrix):

Whut? +_+

OpenStudy (anonymous):

im so lost...

OpenStudy (anonymous):

im so sorry..

OpenStudy (anonymous):

im realllly dumb.. scratch that.

zepdrix (zepdrix):

Try to read what I did with the colors. Match up the colors.

OpenStudy (anonymous):

i put cos(2x-x)=cos2xcosx+sin2xsinx

zepdrix (zepdrix):

Ok great! So the left side of the equation simplified!\[\Large\rm \cos\left(\color{royalblue}{2x}-\color{orangered}{x}\right)=1\]

zepdrix (zepdrix):

So we get cos(x)=1, yes? Not x=1 silly billy +_+

OpenStudy (anonymous):

yeaaa lol my baddd so that means its 0 degrees and 180 degrees and 360 degrees right? then make it into radians?

zepdrix (zepdrix):

No, just 0 and 360. Not 180. Cosine at 180 is -1 yes?

OpenStudy (anonymous):

ohhh yeaa right

zepdrix (zepdrix):

But careful! Look back at your interval, did you post it correctly? Or should one of the sides have square bracket?

zepdrix (zepdrix):

Because \(\Large (0,2\pi)\) gives us no solutions.

OpenStudy (anonymous):

well its 0 degrees greater or equal to x greater or equal to 360 degrees

OpenStudy (anonymous):

no wait less then or equal to

zepdrix (zepdrix):

\[\Large\rm 0\le x\le 360\]Like that?

OpenStudy (anonymous):

yaaa sorr i could only do > and < lol

OpenStudy (anonymous):

i cant make that line under it lol

zepdrix (zepdrix):

In equation tool it's \le short for less than :P Kinda hard to remember all those silly shortcuts though hehe

zepdrix (zepdrix):

Ok if we have the "or equal to" on both equalities, then both sides are square brackets. \(\Large\rm [0,360]\)

zepdrix (zepdrix):

So we include both end points. So yes, 0 and 360.

OpenStudy (anonymous):

so i leave it in degrees right?

zepdrix (zepdrix):

It doesn't matter. Since the problem was given in degrees, yes it's a good idea to leave it in degrees.

OpenStudy (anonymous):

okay THANK YOU JESUS zepdrix thank you thank you thank you i love you

zepdrix (zepdrix):

Lol -_- np.

OpenStudy (anonymous):

man i have a pre calc text tomorrow.. wish me luck

zepdrix (zepdrix):

Ooo fun \c:/ Good luck!

zepdrix (zepdrix):

Trig is the best most fun subject ever :U Don't slack off.

zepdrix (zepdrix):

It's full of amazing identities and great stuff that you'll never ever use unless you decide to become a math major! :D

OpenStudy (anonymous):

thanks. :) like i understand it i just get really unsure about my problem solving so then i get unsure so then i always feel like im doing it wrong

zepdrix (zepdrix):

Mm I'm sure you'll do great c:

OpenStudy (anonymous):

i have a B my grade depends on this test and a final on tuesday.. *gulp*